A father is three times as old as his son. In 14 years' time, he will be twice as old as his son. How old is the father now?
Answer by RAY100(1637)

Let F=fathers age now,,,,and S= son's age now
.
(1) F=3s
.
(2) (F+14) = 2(S+14)
.
F+14 = 2S +28
.
subst from (1),,,,(3S) +14 = 2s +28
.
S= 14,,,,,and F=3(14) =42
.
checking,,,(42 +14) = 2(14) + 28
.
56 = 56,,,,ok
.